pure acid is to be added to a 10% acid solution to obtain 90L of 51% solution. what amounts of each should be used?
how many liters of 100% pure acid should be used to make the solution?
how many liters of the 10% solution should be used in the mixture?

Answer by Alan3354(69443)

pure acid is to be added to a 10% acid solution to obtain 90L of 51% solution. what amounts of each should be used?
how many liters of 100% pure acid should be used to make the solution?
how many liters of the 10% solution should be used in the mixture?
------------------
p + t = 90 (total solution)
1*p = 0.1*t = 0.51*90 (total acid)
-------------
Can you finish it?
-----------
PS Always add the pure acid to the weaker solution, in case it splatters due to heat. Never pour the water, or weaker solution, into strong acid.